A basketball player gets 2 free-throw shots when she is fouled by a player on the opposing team. She misses the first shot 40% of the time. When she misses the first shot, she misses the second shot 5% of the time. What is the probability of missing both free-throw shots?

Verified Answer :

This can easily be answered. Justg use the next formula like this:
0.4*0.05 = 0.02 
So the probability would be of 2% of both shots
Hpe this is very useful

Answer:

\(0.2\) %

Step-by-step explanation:

Chance of missing the first shot is \(0.4\\\)

Chance of missing the second shot, when the first shot is missed is \(0.05\\\)

The probability of missing both free-throw shots

\(= (0.4) (0.05)\\= 0.002\\= 0.2\) %